# Why Do Class-Dependent Evaluation Effects Occur with Time Series Feature Attributions? A Synthetic Data Investigation

### v1

@@ Abstract
Evaluating feature attribution methods represents a critical challenge in explainable AI (XAI), as researchers typically rely on perturbation-based metrics when ground truth is unavailable. However, recent work demonstrates that these evaluation metrics can show different performance across predicted classes within the same dataset. These “class-dependent evaluation effects” raise questions about whether perturbation analysis reliably measures attribution quality, with direct implications for XAI method development and the trustworthiness of evaluation techniques. We investigate under which conditions these class-dependent effects arise by conducting controlled experiments with synthetic time series data where ground truth feature locations are known. We systematically vary feature types and class contrasts across binary classification tasks, then compare perturbation-based degradation scores with ground truth-based precision-recall metrics using multiple attribution methods. Our experiments demonstrate that class-dependent effects emerge with both evaluation approaches even in simple scenarios with temporally localized features, triggered by basic variations in feature amplitude or temporal extent between classes. Most critically, we find that perturbation-based and ground truth metrics frequently yield contradictory assessments of attribution quality across classes, with weak correlations between evaluation approaches. These findings suggest that researchers should interpret perturbation-based metrics with care, as they may not always align with whether attributions correctly identify discriminating features. These findings reveal opportunities to reconsider what attribution evaluation actually measures and to develop more comprehensive evaluation frameworks that capture multiple dimensions of attribution quality. .
@@
However, Baer et al. [1] recently identified a potential methodological limitation: “class-dependent perturbation effects”, where perturbation-based metrics demonstrate different performance across predicted classes within datasets. This phenomenon creates a problem because benchmarks relying on average scores can mask class-specific biases, potentially leading researchers to incorrect conclusions about attribution method quality. Although previous research acknowledges that perturbation results vary by perturbation strategy and recommends testing multiple approaches [16, 13], class-dependent effects seem to persist across various perturbation methods and datasets [1]. The conditions generating these effects remain unexplored, leaving a gap in our understanding of how to reliably evaluate feature attributions.
@@
This paper offers three primary contributions to XAI evaluation methodology. First, we identify specific conditions under which class-dependent evaluation differences emerge, demonstrating that these effects manifest even in minimally complex classification tasks. We show that simple variations in feature prevalence and amplitude across classes trigger class-dependent effects, even when the underlying discriminating feature remains structurally identical between classes. Second, we reveal that class-dependent effects occur in both ground truth-based and perturbation-based evaluation metrics, indicating that attribution methods genuinely perform differently across classes rather than reflecting measurement artifacts. Third, and most critically, we demonstrate that ground truth and perturbation metrics can yield contradictory assessments of attribution quality across classes, revealing that these approaches may measure fundamentally different aspects of attribution behavior. While ground truth metrics assess whether attributions locate discriminating features, perturbation analysis may capture other properties that do not necessarily align with feature identification accuracy. By showing when and why evaluation metrics diverge, our work advances efforts toward more rigorous assessment frameworks in time series analysis as advocated by Theissler et al. [20].
@@ 2 Evaluation Framework
Feature attribution methods for time series classification identify the time points that influence a model’s predictions. Since ground truth is typically unavailable for real-world data, perturbation analysis is commonly used to assess attribution quality by modifying input features and observing the impact on model predictions. The underlying assumption is that perturbing important time points should cause significant prediction changes, while modifying irrelevant time points should have minimal impact. In this work, we evaluate feature attributions using both perturbation-based metrics and ground truth comparison on synthetic data where true feature locations are known. See also Figure 1 for an overview of the evaluation framework.
@@ 2.3 Ground Truth-Based Attribution Evaluation
computed using the trapezoidal rule over unique recall values. We choose AUC-PR because it suits our sparse ground truth features well. For binary classification problems, the random baseline equals the feature prevalence rather than 0.5.
@@ 3 Experimental Design
Class contrasts. We create binary classification tasks by systematically contrasting features across two dimensions: amplitude differences to examine magnitude-based discrimination, and length differences to investigate temporal prevalence effects. This design isolates how different types of class distinctions influence attribution evaluation metrics. We embed features within background Gaussian noise, with feature amplitudes chosen to maintain detectability while avoiding trivial separability. Feature lengths span 20% to 40% of the time series to represent moderate temporal prevalence. Combining four feature types with two contrast mechanisms yields eight datasets. Table 2 summarizes the specific class differences for each contrast type. For example, amplitude-contrast datasets vary only feature magnitude while maintaining identical temporal extent, whereas length-contrast datasets vary temporal coverage while preserving signal strength.
@@
Model architecture and training. We employ two established deep learning architectures for time series classification: ResNet [22] and InceptionTime [6], both of which offer robust performance across time series benchmarks [3]. We train models using AdamW optimization with cosine annealing learning rate scheduling and early stopping (patience = 10 epochs) based on validation performance over a maximum of 100 epochs. Both architectures achieve test accuracy and weighted F1 scores exceeding 0.9 across all datasets, with one exception: InceptionTime on the Amplitude-Pulse dataset (F1 = 0.7). These performance levels ensure that observed class-dependent evaluation differences stem from attribution or perturbation methodology rather than classifier bias toward specific classes. Complete performance metrics are provided in Table 3 (Appendix 0.A).
@@
Perturbation analysis. We employ two perturbation strategies: replacing values with zero and substituting with Gaussian noise. Both strategies are commonly used in time series attribution evaluation [13, 16, 21]. Since we standardize all time series, zero substitution effectively replaces values with the sample mean. For Gaussian noise perturbation, we sample from the same distribution as the background noise (), following the controlled masking approach used in existing benchmarks [5]. This approach ensures that perturbations align more closely with the data distribution, avoiding the potential artifacts of more extreme substitution methods or constant replacement values. We implement perturbation according to MoRF and LeRF orderings determined by attribution scores, substituting features incrementally one timestep at a time. To maintain fair comparisons across datasets with varying feature lengths, we perturb exactly timesteps for each sample, where corresponds to the true feature window size. For instance, in length-contrast datasets, we perturb 30 timesteps for class 0 samples and 60 timesteps for class 1 samples. We record predicted probabilities after each perturbation step to calculate the DS.
@@ 4 Results and Discussion
Ground truth-based evaluation. The ground truth evaluation reveals substantial variation in how well attribution methods recover true class-discriminating feature regions, both across datasets and between classes within datasets. Class 1 observations (with higher amplitude or longer feature windows) consistently achieve higher scores than class 0 observations across seven of eight datasets. The Length-Level dataset is the sole exception where class 1 attributions approach random performance. Overall, Pulse, Sine, and Trend features demonstrate pronounced class-dependent differences in attribution quality. The disparities prove particularly large for Amplitude-Sine and Length-Pulse datasets: class 1 observations in both achieve scores more than double those of class 0 observations, indicating that attribution methods recover ground truth features far more effectively for one class than the other.
@@
Perturbation-based evaluation. Perturbation-based evaluation through DS scores reveals patterns that directly contradict the ground truth findings. Ground truth-based evaluation favors class 1 attributions in seven of eight datasets, while perturbation analysis indicates better attribution quality for class 0 observations in six of eight datasets. This systematic reversal across most feature types and contrast mechanisms suggests the two evaluation approaches measure fundamentally different aspects of attribution behavior. Only two datasets show convergence between evaluation approaches. Both Length-Level and Amplitude-Level datasets exhibit minimal class discrepancies, with both metrics assigning relatively low scores that reflect the sparse attribution patterns characteristic of level shift detection. Only two datasets demonstrate consistent patterns across evaluation approaches. Both Length-Level and Amplitude-Level datasets show smaller class discrepancies within their respective metrics, with both approaches assigning relatively low scores. These lower scores may reflect how attribution methods concentrate on temporal change points rather than extended feature windows when detecting level shifts. This localized attribution behavior could reduce overlap with our ground truth evaluation windows and limit perturbation impact when removing individual timesteps within the feature region. Interestingly, datasets that achieve the highest scores for one class also exhibit the most pronounced disagreement about which class receives superior attribution quality. This pattern appears in both ground truth and perturbation-based evaluation, demonstrating the tension between evaluation approaches even when overall performance appears strong.
@@
Correspondence between evaluation approaches. To examine the relationship between evaluation approaches, we compute Spearman correlations between and DS scores. We use Spearman correlation because the metrics operate on different scales and we focus on monotonic relationships, that is, whether higher scores in one metric correspond to higher scores in the other. Figure 4 shows correlations between both metrics per dataset and class. The analysis confirms our earlier observations about contradictory class-dependent evaluation effects. Correlations remain weak to moderate, ranging from -0.18 to 0.291, and vary substantially between classes within datasets. For example, Amplitude-Sine shows a negative correlation (-0.18) for class 0 but a positive correlation (0.112) for class 1. Similarly, Length-Sine achieves the highest correlation (0.291) for class 1 while class 0 shows a slight negative correlation. These results demonstrate that perturbation-based and ground truth metrics seem to provide different assessments of attribution quality. The weak correspondence indicates that perturbation analysis does not reliably identify whether feature attributions correctly locate true class-discriminating regions in our experimental settings.
@@
Summary. Class-dependent evaluation differences emerge across almost all feature types and contrast mechanisms. Ground truth and perturbation-based metrics yield contradictory assessments: while ground truth evaluation favors class 1 attributions in seven of eight datasets, perturbation analysis indicates better performance for class 0 in six of eight cases. Correlations between approaches remain weak and inconsistent, ranging from -0.18 to 0.291 across datasets and classes.
@@ 5 Conclusion
We investigate conditions under which class-dependent evaluation effects arise in time series feature attribution assessment. Through controlled synthetic experiments with known ground truth features, we compare perturbation-based and ground truth-based evaluation metrics across multiple attribution methods. Our results demonstrate that even simple variations in feature characteristics, such as amplitude differences or temporal extent, can trigger pronounced evaluation disparities between classes. These class-dependent effects emerge consistently across feature types and contrast mechanisms. Most notably, the two evaluation approaches often disagree: perturbation-based metrics frequently suggest superior attribution quality for one class while ground truth metrics indicate the opposite.
These findings highlight important considerations for XAI evaluation in time series domains. While perturbation-based evaluation remains valuable when ground truth is unavailable, our results suggest researchers should interpret such metrics cautiously, recognizing they may not reflect whether attributions correctly identify discriminating features. When synthetic or annotated data permit ground truth comparison, using both evaluation approaches reveals potential contradictions rather than providing validation, as demonstrated in our experimental conditions. We believe that such challenges likely persist and may even compound in more realistic time series datasets.
@@
Our deliberate focus on simplified synthetic scenarios with temporally localized features successfully isolates some of the conditions under which class-dependent effects can emerge, demonstrating that these evaluation challenges arise even under minimal complexity. However, several limitations constrain the generalizability of our findings. We examine only simple univariate time series with temporally localized features, employing two deep learning architectures and three general-purpose attribution methods. While this design enables controlled analysis, it excludes complex temporal patterns and specialized attribution approaches designed for time series. Our evaluation also assumes that good attributions should cover entire feature windows, yet models may legitimately focus on sparse temporal markers like change points. When models identify features through different strategies than we expect, standard metrics may underestimate their actual attribution quality. This mismatch between how we measure attribution performance and how models process temporal information highlights the challenge of defining “correct” attributions in time series contexts.
@@
Future research can address both empirical and conceptual dimensions of these evaluation challenges. Testing time series-specific attribution methods like LimeSegment [18] may reveal whether specialized approaches reduce class-dependent effects. Expanding synthetic datasets to include multiple discriminating patterns and more complex temporal dependencies would test generalizability beyond our localized feature settings. More fundamentally, the weak correspondence between perturbation and ground truth metrics highlights the need to reconsider what attribution evaluation actually measures. For instance, when models correctly identify temporal change points but evaluation metrics expect coverage of entire feature windows, apparent “poor performance” may actually reflect misaligned measurement assumptions. These insights point toward developing evaluation frameworks that embrace multiple perspectives on attribution quality, moving beyond single-metric assessments to better understand how models process temporal information.

### v2

@@ Abstract
Evaluating feature attribution methods represents a critical challenge in explainable AI (XAI), as researchers typically rely on perturbation-based metrics when ground truth is unavailable. However, recent work reveals that these evaluation metrics can show different performance across predicted classes within the same dataset. These “class-dependent evaluation effects” raise questions about whether perturbation analysis reliably measures attribution quality, with direct implications for XAI method development and evaluation trustworthiness. We investigate under which conditions these class-dependent effects arise by conducting controlled experiments with synthetic time series data where ground truth feature locations are known. We systematically vary feature types and class contrasts across binary classification tasks, then compare perturbation-based degradation scores with ground truth-based precision-recall metrics using multiple attribution methods. Our experiments demonstrate that class-dependent effects emerge with both evaluation approaches, even in simple scenarios with temporally localized features, triggered by basic variations in feature amplitude or temporal extent between classes. Most critically, we find that perturbation-based and ground truth metrics frequently yield contradictory assessments of attribution quality across classes, with weak correlations between evaluation approaches. These findings suggest that researchers should interpret perturbation-based metrics with care, as they may not always align with whether attributions correctly identify discriminating features. By showing this disconnect, our work points toward reconsidering what attribution evaluation actually measures and developing more rigorous evaluation methods that capture multiple dimensions of attribution quality. .
@@
However, Baer et al. [1] recently identified a potential methodological limitation: “class-dependent perturbation effects”, where perturbation-based metrics demonstrate different performance across predicted classes within datasets. This phenomenon creates a problem because benchmarks relying on average scores can mask class-specific biases, potentially leading researchers to incorrect conclusions about the quality of attribution methods. Although previous research acknowledges that perturbation results vary by perturbation strategy and recommends testing multiple approaches [16, 13], class-dependent effects seem to persist across various perturbation methods and datasets [1]. The conditions generating these effects remain unexplored, leaving a gap in our understanding of how to reliably evaluate feature attributions.
@@
This paper offers three primary contributions to XAI evaluation methodology. First, we identify specific conditions under which class-dependent evaluation differences emerge, demonstrating that these effects manifest even in minimally complex classification tasks. We show that simple variations in feature prevalence and amplitude across classes trigger class-dependent effects, even when the underlying discriminating feature remains structurally identical between classes. Second, we show that class-dependent effects occur in both ground truth-based and perturbation-based evaluation metrics, indicating that attribution methods genuinely perform differently across classes rather than reflecting measurement artifacts. Third, and most critically, we demonstrate that ground truth and perturbation metrics can yield contradictory assessments of attribution quality across classes, revealing that these approaches may measure fundamentally different aspects of attribution behavior. While ground truth metrics assess whether attributions locate discriminating features, perturbation analysis may capture other properties that do not necessarily align with feature identification accuracy. By showing when and why evaluation metrics diverge, our work advances efforts toward more rigorous assessment frameworks in time series analysis as advocated by Theissler et al. [20].
@@ 2 Evaluation Framework
Feature attribution methods for time series classification identify the time points that influence a model’s predictions. Since ground truth is typically unavailable for real-world data, perturbation analysis is commonly used to assess attribution quality by modifying input features and observing the impact on model predictions. The underlying assumption is that perturbing important time points should cause significant prediction changes, while modifying irrelevant time points should have minimal impact. In this work, we evaluate feature attributions using both perturbation-based metrics and ground truth comparison on synthetic data where true feature locations are known.
Figure 1 illustrates our evaluation framework for systematically investigating class-dependent effects. We generate synthetic time series with known discriminative features, compute attributions using multiple methods, and assess attribution quality through both perturbation-based and ground truth-based evaluations.
Table 1 outlines these complementary approaches, which capture different aspects of attribution quality. Perturbation analysis (Section 2.1) tests whether models rely on attributed regions for predictions, while ground truth evaluation (Section 2.3) verifies whether attributions correctly locate discriminative features. Although these metrics measure distinct properties, we would expect them to agree: attributions that correctly identify discriminative features should also prove critical when perturbed. Our experiments examine when and why these evaluation approaches converge or diverge across classes, showing whether class-dependent patterns reflect genuine attribution differences or evaluation inconsistencies.
| Approach | Metric | Core question and method |
| Perturbation | DS |
Do predictions degrade when removing important features?
Method: Compare model confidence after perturbing most relevant vs. least relevant features. |
| Ground truth | AUC-PR′ |
Do high attribution scores overlap with known discriminative feature regions?
Method: Measure how well attributions retrieve true feature locations using normalized precision-recall analysis. |
@@ 2.3 Ground Truth-Based Attribution Evaluation
computed using the trapezoidal rule over unique recall values. We choose AUC-PR because it suits our sparse ground truth features well. For binary classification problems, the random baseline equals the feature prevalence, .
@@ 3 Experimental Design
Class contrasts. We create binary classification tasks by systematically contrasting features across two dimensions: amplitude differences to examine magnitude-based discrimination, and length differences to investigate temporal prevalence effects. This design isolates how different types of class distinctions influence attribution evaluation metrics.
We embed features within background Gaussian noise, with feature amplitudes chosen to maintain detectability while avoiding trivial separability. Feature lengths span 20% to 40% of the time series to represent moderate temporal prevalence. Combining four feature types with two contrast mechanisms yields eight datasets. Table 3 summarizes the specific class differences for each contrast type. For example, amplitude-contrast datasets vary only feature magnitude while maintaining identical temporal extent, whereas length-contrast datasets vary temporal coverage while preserving signal strength.
Model architecture and training. We employ two established deep learning architectures for time series classification: ResNet [22] and InceptionTime [6], both of which offer robust performance across time series benchmarks [3]. We train models using AdamW optimization with cosine annealing learning rate scheduling and early stopping (patience = 10 epochs) based on validation performance over a maximum of 100 epochs. Both architectures achieve test accuracy and weighted F1 scores exceeding 0.9 across all datasets, with one exception: InceptionTime on the Amplitude-Pulse dataset (F1 = 0.7). These performance levels ensure that observed class-dependent evaluation differences stem from attribution or perturbation methodology rather than classifier bias toward specific classes. Complete performance metrics are provided in Table 4 in Appendix 0.A.
@@
Perturbation analysis. We employ two perturbation strategies: replacing values with zero and substituting with Gaussian noise. Both strategies are commonly used in time series attribution evaluation [13, 16, 21]. Since we standardize all time series, zero substitution effectively replaces values with the sample mean. For Gaussian noise perturbation, we sample from the same distribution as the background noise (), following the controlled masking approach used in existing benchmarks [5]. This approach ensures that perturbations align more closely with the data distribution, avoiding the potential artifacts of more extreme substitution methods or constant replacement values.
We implement perturbations according to MoRF and LeRF orderings determined by attribution scores, substituting features incrementally one timestep at a time. To maintain fair comparisons across datasets with varying feature lengths, we perturb exactly timesteps for each sample, where corresponds to the true feature window size. For instance, in length-contrast datasets, we perturb 30 timesteps for class 0 samples and 60 timesteps for class 1 samples. We record predicted probabilities after each perturbation step to calculate the DS.
@@ 4 Results and Discussion
Ground truth-based evaluation. The ground truth evaluation reveals substantial variation in how well attribution methods recover true class-discriminating feature regions, both across datasets and between classes within datasets. Class 1 observations (with higher amplitude or longer feature windows) consistently achieve higher scores than class 0 observations across seven of eight datasets. The Length-Level dataset is the sole exception where class 1 attributions approach random performance. Overall, Pulse, Sine, and Trend features demonstrate pronounced class-dependent differences in attribution quality. The disparities prove particularly large for Amplitude-Sine and Length-Pulse datasets: class 1 observations in both achieve scores more than double those of class 0 observations, indicating that attribution methods recover ground truth features more effectively for one class than the other.
@@
Perturbation-based evaluation. Perturbation-based evaluation through DS scores reveals patterns that directly contradict the ground truth findings. While ground truth evaluation favors class 1 attributions in seven of eight datasets, perturbation analysis indicates better attribution quality for class 0 observations in six of eight datasets. This systematic reversal across most feature types and contrast mechanisms suggests the two evaluation approaches measure different aspects of attribution behavior. Length-Level and Amplitude-Level datasets represent exceptions where both evaluation approaches converge. These datasets show smaller class discrepancies and assign relatively low scores across both metrics. This pattern likely reflects how attribution methods focus on temporal change points rather than extended feature windows when detecting level shifts. Such localized attribution behavior reduces overlap with our ground truth evaluation windows and limits perturbation impact when removing individual timesteps.
Overall, we observe an inverse relationship: datasets where one class achieves the highest attribution performance also tend to show the strongest disagreement between evaluation approaches. This pattern appears consistently, revealing tensions between ground truth and perturbation-based assessment even when individual metrics indicate strong performance.
Correspondence between evaluation approaches. To examine the relationship between evaluation approaches, we compute Spearman correlations between and DS scores. We use Spearman correlation because the metrics operate on different scales and we focus on monotonic relationships, that is, whether higher scores in one metric correspond to higher scores in the other.
@@
Figure 4 shows correlations per dataset and class. The analysis confirms our earlier observations about contradictory class-dependent evaluation effects. Correlations remain weak to moderate, ranging from -0.18 to 0.291, and vary substantially between classes within datasets. For example, Amplitude-Sine shows a negative correlation (-0.18) for class 0 but a positive correlation (0.112) for class 1. Length-Sine achieves the highest correlation (0.291) for class 1 while class 0 shows a slight negative correlation.
@@
Summary. Class-dependent evaluation differences emerge across almost all feature types and contrast mechanisms. Ground truth and perturbation-based metrics yield contradictory assessments: while ground truth evaluation favors class 1 attributions in seven of eight datasets, perturbation analysis indicates better performance for class 0 in six of eight cases. Spearman correlations between evaluation approaches remain weak and inconsistent, ranging from -0.18 to 0.291 across datasets and classes. This weak correspondence demonstrates that perturbation analysis does not reliably identify whether attributions correctly locate class-discriminating regions, indicating that the two evaluation frameworks measure different aspects of attribution behavior.
@@ 5 Conclusion
We investigate conditions under which class-dependent evaluation effects arise in time series feature attribution assessment. Through controlled synthetic experiments with known ground truth features, we compare perturbation-based and ground truth-based evaluation metrics across multiple attribution methods. Our results demonstrate that even simple variations in feature characteristics, such as amplitude differences or temporal extent, can trigger pronounced evaluation disparities between classes. These class-dependent effects emerge consistently across feature types and contrast mechanisms. Most notably, perturbation-based metrics frequently suggest superior attribution quality for one class while ground truth metrics indicate the opposite.
This systematic disagreement creates a conceptual challenge: in the absence of ground truth, researchers rely on perturbation-based metrics knowing they measure model sensitivity, with the implicit assumption that this serves as a reasonable proxy for feature localization accuracy. Our findings challenge this assumption by demonstrating that model sensitivity and correct feature identification can diverge systematically. The disagreement is particularly concerning since real-world datasets lack ground truth references, leaving perturbation as the primary evaluation option. Researchers should interpret perturbation metrics cautiously, recognizing they may not accurately reflect whether attributions succeed in identifying discriminative feature locations despite our expectations that these properties should align.
@@
Our findings show such disagreements emerge even under minimal complexity. However, several limitations constrain the generalizability of our findings. We examine only simple univariate time series with temporally localized features, employing two deep learning architectures and three general-purpose attribution methods. While this design enables controlled analysis, it excludes complex temporal patterns and specialized attribution approaches designed for time series. Our evaluation also assumes that good attributions should cover entire feature windows, yet models may legitimately focus on sparse temporal markers like change points or learn from subtle patterns beyond our designed features. When models identify features through different strategies than we expect, standard metrics may underestimate their actual attribution quality. This mismatch between how we measure attribution performance and how models process temporal information highlights the challenge of defining “correct” attributions in time series contexts.
@@
Future research can address both empirical and conceptual dimensions of these evaluation challenges. Testing time series-specific attribution methods like LimeSegment [18] may reveal whether specialized approaches reduce class-dependent effects. Expanding synthetic datasets to include multiple discriminating patterns and more complex temporal dependencies would test generalizability beyond our localized feature settings. The weak correspondence between perturbation and ground truth metrics suggests that measuring model sensitivity may not adequately assess feature localization accuracy. This disconnect between what we aim to measure and what perturbation analysis actually captures encourages developing more rigorous evaluation frameworks that acknowledge different dimensions of attribution quality.